OpenStudy (anonymous):

integrate 1/ [sqrt x^2(1+sqrt x)^2] dx

OpenStudy (agreene):

\[\int \frac{1}{\sqrt{x^2} (1+\sqrt x)^2}dx\] Is that really the question?

OpenStudy (anonymous):

yes that it. got it for an assignment

OpenStudy (agreene):

Your professor hates you...

OpenStudy (anonymous):

lol, i think so!

OpenStudy (anonymous):

\[ \int \frac{1}{\sqrt{x^2} (1+\sqrt x)^2}dx = \int \frac{1}{x (1+\sqrt x)^2}dx\]

OpenStudy (anonymous):

i'm attempting to follow how you got your answer @foolformath

OpenStudy (agreene):

start by letting u = sqrt x du = 1/(2 sqrt x) dx That gets us to: \[2\int\frac{1}{u(u+1)^2}\] From there we need to do partial fractions... Also, sqrt x^2 doesnt equal x. That works only if x is a positive number, which we cannot assume

OpenStudy (anonymous):

Indeed agreene is right ...... I am just being a fool :P

OpenStudy (anonymous):

x^2 will always be positive?

OpenStudy (anonymous):

no matter what x is

OpenStudy (agreene):

lol thats a very common mistake... gotta remember those plus/minuses :P So, yeah, you need to do partial fractions, then you'll have to do I think two more substitutions and you'll end up with a big algebraic soup...

OpenStudy (anonymous):

so technically what fool said is correct however you'll get +- and that will just be a disaster

OpenStudy (anonymous):

thanks so much!

OpenStudy (agreene):

@taya http://www.wolframalpha.com/input/?i=1%2F+%5Bsqrt+x%5E2%281%2Bsqrt+x%29%5E2%5D+dx You can click show steps on the right--it gets really messy real quick >.<

OpenStudy (anonymous):

yeah the square roots works as follows . sign : $$ \sqrt[n]{a^n} = a \text{ if $n$ is odd } $$ $$ \sqrt[n]{a^n} = |a| \text{ if $n$ is even } $$

OpenStudy (anonymous):

i tried wolframalpha and got even more confused, but I will have another look @ it

OpenStudy (agreene):

and Outkast... the issue is the slope of the eqn when you assume that. |x| = sqrt x^2 is a true statement if even x = sqrt x^2 iff x is odd

OpenStudy (anonymous):

don't use wolfram alpha for indefinite integrals even simple things loooks messy ...

OpenStudy (agreene):

I just realized I went insane when typing that definition for you outkast... look at fool's he did it correctly.

OpenStudy (anonymous):

ooh, okay... i will look @ fools, thanks

OpenStudy (anonymous):

it's actually a lot easier if you substitute x = u^2 because it reduces to: 2 * integral of ( du/(u(1+u)) )

OpenStudy (anonymous):

= ln|x| - 2ln|1 + sqrt(x)| + C

OpenStudy (anonymous):

ooh, that is so simple that way thank you @james